# Ensembled Direct Multi Step forecasting methodology with comparison on macroeconomic and financial data

### v1

@@ Abstract
Accurate forecasts of macroeconomic and financial data, such as GDP, CPI, unemployment rates, and stock indices, are crucial for the success of countries, businesses, and investors, resulting in a constant demand for reliable forecasting models. This research introduces a novel methodology for time series forecasting that combines Ensemble technique with a Direct Multi-Step (DMS) forecasting procedure. This Ensembled Direct Multi-Step (EDMS) approach not only leverages the strengths of both techniques but also capitalizes on their synergy. The ensemble models were selected based on performance, complexity, and computational resource requirements, encompassing a full spectrum of model complexities, from simple Linear and Polynomial Regression to medium-complexity ETS and complex LSTM models. Ensembling is carried out using weights derived from each model’s performance. The DMS procedure limits retraining to one- and five-year forecasts for economic data and one- and five-month forecasts for financial data. The standard Iterative Multi-Step (IMS) procedure is employed for other horizons, effectively reducing computational demands while maintaining satisfactory results. The proposed methodology is benchmarked against the Ensemble technique conventionally applied to IMS-generated forecasts, utilizing several publicly available macroeconomic datasets, including GDP, CPI, and employment figures across selected countries, and common financial indices data. Results demonstrate a significant performance improvement with the EDMS methodology, averaging a 33.32% enhancement across the analysed datasets, and sometimes reaching improvement above 60%.
@@ 1 Introduction
In recent years, time series forecasting has gained significant importance across various fields, particularly in economics, finance, and meteorology. This is primarily due to its crucial role in predicting future trends, which enables better decision-making and risk mitigation. For instance, accurate forecasts of key macroeconomic indicators, such as GDP, CPI, interest rates, and unemployment are essential for central banks, monetary authorities, and governments to guide monetary policy decisions, evaluate economic performance and labor market conditions, and plan government spending and taxation. Similarly, predicting future traffic flow or energy consumption facilitates improved planning in the economy, ensuring optimal resource allocation. Economic forecasting specifically, faces several challenges, including model uncertainty, non-stationarity, and the mixed frequency of data. In finance, forecasts are fundamental for establishing appropriate trading strategies, including hedging, speculation, and arbitrage. Private and corporate investors, businesses, commercial banks, and other financial institutions compute various forecasts daily, covering prices of stocks, indices, derivatives, and other financial assets. Lastly, in meteorology, forecasting weather conditions, including wind strength and weather anomalies is crucial for maritime transport and for protecting populations from unforeseen weather events.
@@
Traditionally, forecasts have been computed as a one-step-ahead prediction problem, aiming to predict the value at the next time step. Longer-term forecasts are generated through the iterative application of the same model across consecutive time steps, a procedure commonly referred to as Iterative Multi-Step (IMS). When the model is retrained at each step by incorporating previously forecasted data, the approach is known as Direct Multi-Step (DMS).
@@
Although DMS is less common than IMS, it offers significant advantages. Most importantly, DMS tends to enhance forecasts, enabling more precise macroeconomic forecasting and planning. According to [14] DMS is most beneficial when the model is misspecified, particularly when the data contains misspecified unit roots, neglected residual autocorrelation, and omitted location shifts. Retraining the model at each step in the forecast horizon generally makes the approach more specialized and tailored for specific use cases. Specifically, at each retraining step, the training set is extended with predictions from previous forecast executions, resulting in forecasts that are naturally more accurate than those generated by IMS. This improved accuracy is especially evident when dealing with non-stationary data or when model dynamics change over time [45, 47, 15]. However, the downside of this approach is the significant consumption of time and computing resources, as retraining is typically the most computationally demanding part of the forecasting process. Moreover, in DMS, we can utilize different models for each forecast step, which tend to perform better with longer forecast horizons [14, 45, 68]. Nonetheless, this approach presents additional challenges, such as reliance on varying sets of input parameters, model structures, or even learning algorithms, which adds complexity to the setup and execution of the training process [68].
@@
In contrast, the traditional IMS method involves creating a model tailored to a specific date within the forecasting horizon and using it repeatedly for consecutive time steps. This approach simplifies the forecasting process and, in some cases, can achieve similar accuracy to DMS, particularly when the data is stationary and the model is well-tuned [14, 45]. Notably, iterated forecasting demonstrates better performance in terms of computing resources compared to DMS, as the model is trained only once. This is especially beneficial when dealing with numerous time series and long forecast horizons. However, as discussed in [68], the IMS approach can struggle to fit the data adequately, even when the model is perfectly specified for the data-generating process. Moreover, IMS is more sensitive to model misspecification than DMS, as shown in [68], the Mean Squared Error (MSE) of IMS tends to be larger than that of DMS if the model is incorrectly specified.
@@
DMS can be preferred over IMS, especially for macroeconomic forecasting, due to several practical and theoretical advantages. In IMS, each forecasted value serves as an input for the subsequent forecast using the same model. This can lead to the accumulation of errors. A small inaccuracy in earlier predictions can accumulate over time, resulting in significantly less accurate forecasts for later periods. In contrast, DMS allows each model to be optimized for its target period, often leading to better performance, particularly in the context of macroeconomic indicators such as GDP, inflation, or employment. These indicators are influenced by numerous factors that DMS can capture more effectively. Additionally, economic indicators frequently exhibit non-stationary behaviour. DMS forecasting can better accommodate these changes, as the model can be designed to incorporate recent trends and shifts in the data directly. Furthermore, macroeconomic data often display nonlinear behavior and are subject to structural breaks due to policy changes, economic shocks, or other external factors. DMS forecasting models are better equipped to handle these complexities because they do not rely on recursively generated inputs that may carry forward outdated or biased information.
@@
A common technique for improving forecast accuracy is Ensembling. In this approach, an ensemble of models is utilized to compute forecasts rather than relying on a single model. The ensemble forecast is calculated as a weighted sum of the forecasts produced by individual models. The weights can either be equal or determined based on each model’s performance. Research indicates that ensemble forecasting tends to generalize better to unseen data, as it leverages the strengths of multiple models while complementing their weaknesses. Consequently, the ensembled forecast is typically more accurate than those generated by individual models.
@@
We propose a new forecasting methodology called Ensembled Direct Multi-Step (EDMS), which combines the Direct Multi-Step approach with the Ensemble forecasting technique. Similar to the Direct Multi-Step (DMS) method, EDMS involves retraining models. However, we limit retraining to specified time steps, during which forecasts are generated using the Ensembling technique. For the remaining time steps, forecasts are iteratively computed, as in the Iterative Multi-Step (IMS) method. In general, retraining in EDMS can be performed at arbitrary time steps throughout the forecasting horizon. Reducing the number of retrainings is particularly important when forecasts consume significant computing resources. This situation often arises when dealing with large training datasets, such as an extensive number of time series or observations, which is common in economic and financial datasets. The models included in the ensemble are carefully selected based on criteria such as accuracy, complexity, and computational efficiency. When ensembling, each model contributes to the ensemble according to its weight, which reflects its performance. This ensures that the most accurate forecasts have the greatest impact on the ensemble. To assess the practical viability of EDMS, we conduct an extensive comparative analysis of the EDMS methodology against the conventional Ensemble technique that utilizes IMS (EIMS). This analysis is performed on common macroeconomic and financial datasets.
@@
The remainder of the paper is organized into several sections: Section 2 offers a comprehensive literature review of the methodology’s components, specifically the Ensembled Forecasting Method, the models utilized in the ensemble, and the IMS and DMS procedures, as well as some improvements of DMS. Section 3 details the EDMS methodology, including a description of the models employed in the ensemble and the Ensemble technique itself. In Section 4, the document presents a comparative analysis of macroeconomic data, evaluating the performance of the EDMS methodology in relation to the conventional Iterative Multi-Step forecasting approach. This section also characterizes the datasets used and the comparison methodology, including performance metrics such as the Mean Absolute Percentage Error (MAPE). At the end of Section 4, the results are presented, showing significant improvements in forecasting accuracy with the use of the EDMS approach. The final section concludes with a summary of the key findings of the study, as well as potential improvements, extensions and applications.
@@ 2 Literature review
In the literature, there is a wide variety of techniques available for forecasting time series. Recently, much of the research has focused on machine learning (ML) models, particularly Artificial Neural Networks (ANNs). While traditional models, such as ARIMA and ETS, have been successfully used for years, the application of ANNs offers significant improvements in capturing complex temporal patterns. Notably, the most researched ANN models are Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRU), which have been well-established to produce accurate forecasts [52, 41, 29, 64, 56, 11, 66, 42]. Furthermore, LSTM and GRU are often utilized as components of hybrid approaches, such as in combination with ARIMA. Several articles demonstrate that such hybrid approaches perform better than the individual models alone [38,39,40,41].
@@
Furthermore, when it comes to forecasting economic time series, machine learning models frequently outperform traditional econometric models, such as VAR or (S)ARIMA(X). In particular, artificial neural network (ANN) architectures, especially LSTM and GRU, have proven to be well-suited for this purpose [67, 2, 22, 40]. Additionally, the literature presents numerous comparisons of a wide range of machine learning models, including XGBoost, LSTM, GRU, and DeepAR, against traditional ARIMA. In almost all cases, machine learning models demonstrate superior performance compared to the traditional counterpart [4, 24, 33, 77, 38]. One explanation for this phenomenon is the high dependence of machine learning on hyperparameter selection, which allows for extensive tuning and offers the potential to significantly enhance accuracy.
@@
Numerous comparisons have been made between the well-known Prophet model [1] and both machine learning (ML) and traditional econometric models. It is important to note that Prophet is an enhancement of traditional models. Specifically, Prophet generalizes the additive model, where different characteristics of the time series, such as trend, seasonality, holidays, and special events, are combined, along with the addition of random noise [1]. Some studies assert that Prophet outperforms traditional models [13, 21, 3, 39, 10, 63], while others suggest that ML or hybrid ML models are superior [4, 24, 33, 77, 38, 25]. The reason behind Prophet’s strong performance lies in its specific design, which is also particularly well-suited for handling seasonal data effectively.
@@
Ensemble forecasting is a widely adopted strategy for forecasting. This technique aggregates predictions from multiple models. Such approaches consistently outperform individual models in terms of accuracy and stability across diverse domains. The seminal work of [6] first demonstrated the benefits of combining forecasts, highlighting significant improvements in predictive performance. Subsequent studies by [49, 75] confirmed the effectiveness of both linear and nonlinear combinations. Building on this foundation, [36, 35] showed that combining models trained on data sampled at different frequencies enhances forecasting accuracy across short-, medium-, and long-term horizons, with minimal differences between mean and median operators. Additional empirical evidence from studies such as [76, 8, 55, 12, 26, 53, 37, 5, 60, 70, 57, 51, 9, 65] and [16], p. 145, further shows the effectiveness of ensemble methods in improving forecast performance.
@@
The concept of Direct Multi-step forecasting procedure has been around for a relatively long time, and the debate between IMS and DMS remains active, with various studies comparing their performance. Many research studies indicate that DMS outperforms IMS, however, there are also examples demonstrating the opposite. According to the work of Chevillion and Hendry [15], early research conducted by Klein [34] and Johnston, Klein, and Shinjo [30] suggested that DMS might be more efficient than IMS. These results were inspired by the application of Cox’s [18] ideas to exponentially-weighted moving-average (EWMA) or integrated moving-average (IMA (1,1)) models [15]. In extensive empirical studies, Kang [32] and Marcellino et al. [46] observed only minimal improvements with DMS, while Proietti [54] extended these findings to ARIMA models and reached a similar conclusion. According to the works of Bhansali [7] and Schorfheide [61], under certain assumptions, DMS can outperform IMS. Bhansali asserts that for a finite autoregressive process, IMS is superior; however, for other models, DMS can effectively reduce the mean squared error of prediction when using an under-parametrized model. In contrast, Schorfheide demonstrates that DMS forecasting yields favorable results for series related to wages, prices, and money. Furtehmore, Chevillon (2017) highlights that DMS models are particularly effective when managing the bias-variance trade-off over extended forecasting horizons. In contrast, more recently, McCracken et al. [47] compared the accuracy of direct multi-step and iterated multi-step forecasting using VAR and ARDL models, specifically VAR-based IMS and ARDL-based DMS. They concluded that IMS approaches generally perform better; however, DMS-based methods show promise for forecasting hypothetical scenarios, such as the influence of oil prices on inflation in a VAR model and for nominal variables.
Numerus study shows DMS tends to give better results when the used model is misspecifies underlying data generating process of the modelled phenomoena. Moreover, Lin and Tsay [43] demonstrated that certain types of model misspecification can make better forecast when we use DMS instead of IMS [43, 71, 32, 46, 15]. With similar conclusion came in their paper Tiao and Xu’s [69], that is, DMS is more efficient than IMS if the model is mis-specified. On the other hand, when the model is correctly specified, Johnston [30] indicate that IMS performs better [18, 34, 15]. In contrary, in another study, Weiss [73] explored DMS and ARIMA models. He’s results indicate that when the model is incorrectly specified, IMS and DMS gave relatively good results, although in small samples, that is, when the number of observations is limited, DMS works worse [20, 73, 74, 15]. Further claims around misspecification provides Chevillon [14]. He concludes that DMS is a valuable technique for forecasting, especially in the presence of model misspecification and non-stationarity. He also finds that the benefits of IMS or DMS vary with the forecast horizon and the stochastic properties of the data.
@@
When comparing DMS to IMS, McElroy [48] argues that the differences in performance between direct and iterative forecasting primarily stem from the model fitting method and the subtleties of model misspecification. Therefore, the choice of forecasting method should be made with careful consideration of the model’s specifics and the context of its application. Numerous studies indicate that DMS tends to yield better results when the employed model misrepresents the underlying data-generating process of the modeled phenomena. Moreover, Lin and Tsay [43] demonstrated that certain types of model misspecification can lead to improved forecasts when using DMS instead of IMS [71, 32, 46, 15]. A similar conclusion was reached in Tiao and Xu’s [69] paper, which states that DMS is more efficient than IMS when the model is mis-specified. Conversely, when the model is correctly specified, Johnston [30] indicated that IMS performs better [18, 34, 30, 15]. In contrast, Weiss [73] explored DMS and ARIMA models, finding that when the model is incorrectly specified, both IMS and DMS yield relatively good results. However, in small samples, when the number of observations is limited, DMS tends to perform worse [20, 73, 74, 15]. Further insights regarding misspecification are provided by Chevillon [14], who concludes that DMS is a valuable technique for forecasting, particularly in the presence of model misspecification and non-stationarity. He also found that the advantages of IMS or DMS vary depending on the forecast horizon and the stochastic properties of the data.
@@
Additionally, literature provide us with some improvements of DMS procedure. For instance, In and Jung [28] propose a mixed approach of DMS with IMS, known as the Direct and Recursive Forecast Averaging Method (DRFAM). This method integrates various strategies for forecasting daily product sales over the next 28 days using hierarchical data. The features of both DMS and IMS are essential to the DRFAM model, and their combined application improves forecasting accuracy when addressing challenges associated with hierarchical data. Furthermore, Dossani [19] introduces a novel method for inference in direct multi-step forecasting regressions. This approach involves estimating the covariance matrix in direct multi-step and long-horizon forecasting regressions, which enhances both inference accuracy and efficiency. The proposed estimator significantly improves efficiency and accuracy compared to conventional HAC methods (heteroskedasticity and autocorrelation consistent covariance estimator) for both direct multi-step and long-horizon forecasts. The inference method presented in this paper enhances forecasting by providing more accurate estimates of uncertainty (standard errors) around forecasted values, thereby increasing confidence in predictive models. Clark and West [17] investigate the performance of DMS forecasting in nested models, with a particular focus on utilizing mean squared prediction error (MSPE). Their paper compares two nested linear forecasting models: a parsimonious model with fewer parameters and a model that includes additional predictors. The authors evaluate the out-of-sample predictive performance of these models using adjusted mean squared prediction error (MSPE). The final empirical applications demonstrate that the MSPE-adjusted measure provides a reliable assessment of forecast accuracy in nested models, underscoring its practical relevance in forecasting Econometricians continue to explore both, direct and iterated, multi-step forecasting methods, with significant contributions from Findley, Weiss, Tiao, Xu, Lin, Granger, Clements, Hendry, Bhansali, Kang, Chevillon, and Schorfheide.

### v2

# Ensembled Direct Multi-Step forecasting methodology with comparison on macroeconomic and financial data

@@ Abstract
Accurate forecasts of macroeconomic and financial data, such as GDP, CPI, unemployment rates, and stock indices, are crucial for the success of countries, businesses, and investors, resulting in a constant demand for reliable forecasting models. This research introduces a novel methodology for time series forecasting that combines Ensemble technique with a Direct Multi-Step (DMS) forecasting procedure. This Ensembled Direct Multi-Step (EDMS) approach not only leverages the strengths of both techniques but also capitalizes on their synergy. The ensemble models in the methodology were selected based on performance, complexity, and computational resource requirements, encompassing a full spectrum of model complexities, from simple Linear and Polynomial Regression to medium-complexity ETS and complex LSTM model. Models were weighted based on their performances. In the DMS procedure we limit retraining to one- and five- year/month forecasts for economic and financial data respectively. Standard Iterative Multi-Step (IMS) procedure is employed for other horizons, effectively reducing computational demands while maintaining satisfactory results. The proposed methodology is benchmarked against Ensemble technique conventionally applied to IMS-generated forecasts, utilizing several publicly available macroeconomic and financial datasets. Results demonstrate a significant performance improvement with EDMS methodology, averaging a 33.32% enhancement across the analysed datasets, and sometimes improvement reaching above 60%.
@@ 1 Introduction
In recent years, time series forecasting has gained significant importance across various fields, particularly in economics, finance, and meteorology. This is primarily due to its crucial role in predicting future trends, which enables better decision-making and risk mitigation. For instance, accurate forecasts of key macroeconomic indicators, such as GDP, CPI, interest rates, and unemployment are essential for central banks, monetary authorities, and governments to guide monetary policy decisions, evaluate economic performance and labour market conditions, and plan government spending and taxation. Similarly, predicting future traffic flow or energy consumption facilitates improved planning in the economy, ensuring optimal resource allocation. Economic forecasting specifically, faces several challenges, including model uncertainty, non-stationarity, and the mixed frequency of data. In finance, forecasts are fundamental for establishing appropriate trading strategies, including hedging, speculation, and arbitrage. Private and corporate investors, businesses, commercial banks, and other financial institutions compute various forecasts daily, covering prices of stocks, indices, derivatives, and other financial assets. Lastly, in meteorology, forecasting weather conditions, including wind strength and weather anomalies is crucial for maritime transport and for protecting populations from unforeseen weather events.
@@
Traditionally, forecasts have been computed as a one-step-ahead prediction problem, aiming to predict the value at the next time step, and longer-term forecasts are generated through the iterative application of the same model across consecutive time steps. This procedure is commonly referred as Iterative Multi-Step (IMS). The refinement of this procedure is when the model is retrained at each step by incorporating previously forecasted data, the approach is known as Direct Multi-Step (DMS).
@@
Although DMS is less common than IMS, it offers significant advantages. Most importantly, DMS tends to enhance forecasts accuracy, enabling more precise macroeconomic forecasting and planning. According to [14] DMS is most beneficial when the model is misspecified, particularly when the data contains misspecified unit roots, neglected residual autocorrelation, and omitted location shifts. Retraining the model at each step in the forecast horizon generally makes the approach more specialized and tailored for specific use cases. Specifically, at each retraining step, the training set is extended with predictions from previous forecast executions, resulting in forecasts that are naturally more accurate than those generated by IMS. This improved accuracy is especially evident when dealing with non-stationary data or when model dynamics change over time [45, 47, 15]. However, the downside of this approach is the significant consumption of time and computing resources, as training is typically the most computationally demanding part of the forecasting process. Moreover, in DMS, we can utilize different models for each forecast step, which tend to perform better with longer forecast horizons [14, 45, 68]. Nonetheless, this approach presents additional challenges, such as reliance on varying sets of input parameters, model structures, or even learning algorithms, which adds complexity to the setup and execution of the training process [68].
@@
In contrast, the traditional IMS method involves creating a model tailored to a specific date within the forecasting horizon and uses it repeatedly for consecutive time steps. This approach simplifies the forecasting process and, in some cases, can achieve similar accuracy to DMS, particularly when the data is stationary and the model is well-tuned [14, 45]. Notably, iterated forecasting demonstrates better performance in terms of computing resources compared to DMS, as the model is trained only once. This is especially beneficial when dealing with numerous time series and long forecast horizons. However, as discussed in [68], the IMS approach can struggle to fit the data adequately, even when the model is perfectly specified for the training data. Moreover, IMS is more sensitive to model misspecification than DMS, as shown in [68], the Mean Squared Error (MSE) of IMS tends to be larger than that of DMS if the model is incorrectly specified.
@@
DMS can be preferred over IMS, especially for macroeconomic forecasting, due to several practical and theoretical advantages. In IMS, the fact that forecasted value serves as an input for the subsequent forecast using the same model, leads to the accumulation of errors. A small inaccuracy in earlier predictions can accumulate over time, resulting in significantly less accurate forecasts for later periods. In contrast, DMS allows each model to be optimized for its target period, often leading to better performance, particularly in the context of macroeconomic indicators such as GDP, inflation, or employment. These indicators are influenced by numerous dynamic factors and frequently exhibit non-stationary behaviour. DMS forecasting can better accommodate these changes, as the model can be designed to incorporate recent trends and shifts in the data more dynamically. Furthermore, macroeconomic data often display nonlinear behaviour and are subject to structural breaks due to policy changes, economic shocks, or other external factors. DMS forecasting models are better equipped to handle these complexities because they do not rely on recursively generated inputs that may carry forward outdated or biased information.
@@
A common technique for improving forecast accuracy is Ensembling. In this approach, an ensemble of models is utilized to compute forecasts rather than relying on a single model. The ensemble forecast is calculated as a weighted sum of the forecasts produced by individual models. The weights can either be equal or determined based on each model’s performance. Research indicates that ensembled forecasts tends to generalize better to unseen data, as this technique leverages the strengths of multiple models while complementing their weaknesses. Consequently, the ensembled forecast is typically more accurate than those generated by individual models.
@@
We propose a new forecasting methodology called Ensembled Direct Multi-Step (EDMS), which combines Direct Multi-Step approach with Ensemble forecasting technique. Similar to DMS method, EDMS involves retraining models. However, we limit retraining to specified time steps, and also for this time steps forecasts are generated using the Ensemble technique. For the remaining time steps, forecasts are iteratively computed, as in IMS method. In general, retraining in EDMS can be performed at arbitrary time steps throughout the forecasting horizon. Reducing the number of retrainings is particularly important when forecasts consume significant computing resources. This situation often arises when dealing with large training datasets, such as an extensive number of time series or observations, which is common in economic and financial datasets. The models included in the ensemble are carefully selected based on criteria such as accuracy, complexity, and computational efficiency. When ensembling, each model contributes to the ensemble according to its weight, which reflects its performance. This ensures that the most accurate forecasts have the greatest impact on the ensemble. To assess the practical viability of EDMS, we conduct an extensive comparative analysis of this methodology against the conventional Ensemble technique that utilizes IMS (EIMS). This analysis is conducted on common and publicly available macroeconomic and financial datasets.
@@
The remainder of the paper is organized into several sections: Section 2 offers a comprehensive literature review of the methodology’s components, specifically Ensembled forecasting method, models utilized in the ensemble, and IMS and DMS procedures, as well as some improvements of DMS. Section 3 details the EDMS methodology, including a description of the models employed in the ensemble and the Ensemble technique itself. In Section 4, the document presents a comparative analysis of macroeconomic data, evaluating the performance of the EDMS methodology in relation to the conventional Iterative Multi-Step forecasting approach. This section also characterizes the datasets used and the comparison methodology. At the end of Section 4, the results are presented, showing significant improvements in forecasting accuracy with the use of the EDMS approach. The final, fifth section, concludes with a summary of the key findings of the study, as well as potential improvements, extensions and applications.
@@ 2 Literature review
In the literature, there is a wide variety of techniques available for forecasting time series. Recently, much of the research focus is on machine learning (ML) models, particularly Artificial Neural Networks (ANNs). While traditional models, such as ARIMA and ETS, have been successfully used for years, the application of ANNs offers significant improvements in many aspects of forecasting, including capturing complex temporal patterns. Notably, the most researched ANN model is Long Short-Term Memory (LSTM), which has been well-established to produce accurate forecasts [52, 41, 29, 64, 56, 11, 66, 42]. Furthermore, LSTM is often utilized as components of hybrid approaches, such as in combination with ARIMA. Several articles demonstrate that such hybrid approaches perform better than the individual models alone [38,39,40,41].
@@
Furthermore, when it comes to forecasting economic time series, machine learning models frequently outperform traditional econometric models, such as VAR or (S)ARIMA(X). In particular, ANN architectures, especially LSTM, has proven to be well-suited for this purpose [67, 2, 22, 40]. Additionally, the literature presents numerous comparisons of a wide range of machine learning models, including XGBoost, LSTM, GRU, and DeepAR, against traditional ARIMA. In almost all cases, ML models demonstrate superior performance compared to the traditional counterpart [4, 24, 33, 77, 38]. One explanation for this phenomenon is variety of available models, so that the model structure and architecture can be tailor to specific use case. Moreover, high dependence of machine learning on hyperparameters selection allows for extensive tuning and offers the potential to significantly enhance accuracy.
@@
Numerous comparisons have been made between the well-known Prophet model [1] and both, ML and traditional econometric models. It is important to note that Prophet is an enhancement of traditional models. Specifically, Prophet generalizes the additive model, where different characteristics of the time series, such as trend, seasonality, holidays, and special events, are combined, along with the addition of random noise [1]. Some studies assert that Prophet outperforms traditional models [13, 21, 3, 39, 10, 63], while others suggest that ML or hybrid ML models are superior [4, 24, 33, 77, 38, 25]. The reason behind Prophet’s strong performance lies in its specific design, which is also particularly well-suited for handling seasonal data effectively.
@@
Ensemble forecasting is a widely adopted strategy for forecasting. This technique aggregates predictions from multiple models. With such a approach forecasts consistently outperform individual models forecasts in terms of accuracy and stability across diverse domains. The seminal work of [6] first demonstrated the benefits of combining forecasts, highlighting significant improvements in predictive performance. Subsequent studies by [49, 75] confirmed the effectiveness of both linear and nonlinear combinations. Building on this foundation, [36, 35] showed that combining models trained on data sampled at different frequencies enhances forecasting accuracy across short-, medium-, and long-term horizons, with minimal differences between mean and median operators. Additional empirical evidence from studies such as [76, 8, 55, 12, 26, 53, 37, 5, 60, 70, 57, 51, 9, 65] and [16], p. 145, further shows the effectiveness of ensemble methods in improving forecast performance.
@@
The concept of Direct Multi-step forecasting procedure has been around for a relatively long time, and the debate on IMS and DMS remains active, with various studies comparing their performance. Many research studies indicate that DMS outperforms IMS, however, there are also examples demonstrating the opposite. According to the work of Chevillion and Hendry [15], early research conducted by Klein [34] and Johnston, Klein, and Shinjo [30] suggests that DMS might be more efficient than IMS. These results were inspired by the application of Cox’s [18] ideas to exponentially-weighted moving-average (EWMA) or integrated moving-average (IMA (1,1)) models [15]. In extensive empirical studies, Kang [32] and Marcellino et al. [46] observed only minimal improvements with DMS, while Proietti [54] extended these findings to other models, i.e. ARIMA, and reached a similar conclusion. According to works of Bhansali [7] and Schorfheide [61], under certain assumptions, DMS can outperform IMS. Bhansali asserts that for a finite autoregressive process, IMS is superior; however, for other models, DMS can effectively reduce the mean squared error of prediction when using an under-parametrized model. In contrast, Schorfheide demonstrates that DMS forecasting yields favourable results for series related to wages, prices, and money. Furthermore, Chevillon (2017) highlights that DMS models are particularly effective when managing the bias-variance trade-off over extended forecasting horizons. In contrast, more recently, McCracken et al. [47] compared the accuracy of direct multi-step and iterated multi-step forecasting using VAR and ARDL models, specifically VAR-based IMS and ARDL-based DMS. They concluded that IMS approaches generally perform better, however, DMS-based methods show promise for forecasting hypothetical scenarios, such as the influence of oil prices on inflation in a VAR model and for nominal variables.
@@
Numerous studies indicate that DMS tends to yield better results when the employed model misrepresents the underlying data-generating process of the modelled phenomena. Moreover, Lin and Tsay [43] demonstrated that certain types of model misspecification can lead to improved forecasts when using DMS instead of IMS [71, 32, 46, 15]. A similar conclusion was reached in Tiao and Xu’s [69] paper, which states that DMS is more efficient than IMS when the model is misspecified. Conversely, when the model is correctly specified, Johnston [30] indicated that IMS performs better [18, 34, 30, 15]. In contrast, Weiss [73] explored DMS and ARIMA models, finding that when the model is incorrectly specified, both IMS and DMS yield relatively good results. However, in small samples, when the number of observations is limited, DMS tends to perform worse [20, 73, 74, 15]. Further insights regarding misspecification are provided by Chevillon [14], who concludes that DMS is a valuable technique for forecasting, particularly in the presence of model misspecification and non-stationarity. He also found that the advantages of IMS or DMS vary depending on the forecast horizon and the stochastic properties of the data. When comparing DMS to IMS, McElroy [48] argues that the differences in performance between direct and iterative forecasting primarily stem from the model fitting method and the subtleties of model misspecification. Therefore, the choice of forecasting method should be made with careful consideration of the model’s specifics and the context of its application.
@@
Additionally, literature provide us with some improvements of DMS procedure. For instance, In and Jung [28] propose a mixed approach of DMS with IMS, known as the Direct and Recursive Forecast Averaging Method (DRFAM). This method integrates various strategies for forecasting daily product sales over the next 28 days using hierarchical data. The features of both DMS and IMS are essential to the DRFAM model, and their combined application improves forecasting accuracy when addressing challenges associated with hierarchical data. Furthermore, Dossani [19] introduces a novel method for inference in direct multi-step forecasting regressions. This approach involves estimating the covariance matrix in direct multi-step and long-horizon forecasting regressions, which enhances both inference accuracy and efficiency. The proposed estimator significantly improves efficiency and accuracy compared to conventional HAC methods (heteroskedasticity and autocorrelation consistent covariance estimator) for both direct multi-step and long-horizon forecasts. The inference method presented in that paper enhances forecasting by providing more accurate estimates of uncertainty, i.e. standard errors, around forecasted values, thereby increasing confidence in predictive models. Clark and West [17] investigate the performance of DMS forecasting in nested models, with a particular focus on utilizing mean squared prediction error (MSPE). Their paper compares two nested linear forecasting models: a parsimonious model with fewer parameters and a model that includes additional predictors. The authors evaluate the out-of-sample predictive performance of these models using adjusted mean squared prediction error (MSPE). The final empirical applications demonstrate that MSPE-adjusted measure provides a reliable assessment of forecast accuracy in nested models, underscoring its practical relevance in forecasting. Econometricians continue to explore both, direct and iterated, multi-step forecasting methods, with significant contributions from Findley, Weiss, Tiao, Xu, Lin, Granger, Clements, Hendry, Bhansali, Kang, Chevillon, and Schorfheide.